# A Quality Diversity Method to Automatically Generate Multi-Agent Path Finding Benchmark Maps

### v3

@@ 4 Benchmarking One Algorithm
For one-algorithm experiments, we intend to generate a diverse set of difficult benchmark maps targeted for five representative algorithms we choose from Section 2.1, namely CBS, EECBS, PBS, PIBT, and LTF.
@@
Figure 2: Examples of one-entry space and one-tile entry, circled in red.
@@ 4.1 Experiment Setup
Hyperparameters. Defining $w$ as the suboptimality bound away from the optimal solution, we use $w=1.5$ for EECBS. For CBS, we use the implementation of EECBS with $w=1$. We set $T=20$ seconds for CBS, EECBS, and PBS, $M=1000$ for PIBT, and $M=512$ for LTF. For all algorithms, we fix the map size as 32 $\times$ 32 and the number of agents at 150. The pre-defined range for the number of obstacles is $[O_{lb}=307,O_{ub}=717]$, which corresponds to 30% to 70% of the map size. For each evaluation, we run the given MAPF algorithm in $N_{e}=5$ different instances. For all algorithms, we evaluated $N_{eval}=10,000$ maps. For CBS, EECBS, and PBS, we run evaluations with $N_{a}=50$ agents, and for PIBT and LTF, we use $N_{a}=150$. We use more agents for PIBT and LTF because they empirically have higher success rates in congested MAPF instances.
Implementation. We implement CMA-MAE with Pyribs (Tjanaka et al. 2023). For MAPF algorithms, we use the open source implementations of CBS and EECBS [^3] https://github.com/Jiaoyang-Li/EECBS, PBS [^4] https://github.com/Jiaoyang-Li/PBS, PIBT [^5] https://github.com/Kei18/pibt2, and LTF [^6] https://github.com/AIRI-Institute/learn-to-follow.
@@ 4.2 Results
We first define one-entry space and one-tile entry (shown in Figure 2) as features of the benchmark maps. In Figures 3, 4 and 5, the yellow cells indicate maps where the algorithm runs out of time. The dark blue cells indicate maps where the algorithm quickly finds a solution. In Figures 6 and 7, the yellow cells indicate maps where the algorithm fails to plan for any agents. The dark blue cells indicate maps where the algorithm successfully plans for all agents. The colored tiles in the representative maps display 10 out of 150 pairs of start and goal locations. While we are primarily interested in maps that are challenging for the algorithms, we also find maps that are easy, as the by-product of the experiment. We discuss more in Section 6.
@@
CBS. Figure 3 shows the archive of maps for CBS. It is easy to generate challenging maps for CBS. Map (a), Map (c), and Map (d) show the most typical patterns where it is hard for CBS to find optimal solutions. Map (a) and Map (c) contain many long corridors, and Map (d) contains many one-entry spaces with long corridors, where one-entry spaces represent the empty space surrounded by three obstacles. Figure 2(a) shows an example. By running CBS on Map (a) and Map (b) in 200 MAPF instances, the success rate is below 5% with an average runtime of over 18 seconds. The result shows that CBS performs poorly on maps with many long corridors and one-entry spaces.
Some maps that CBS can solve are maps with fewer obstacles and a large chunk of empty spaces. Map (b) is one of these typical maps. Map (b) contains a large chunk of empty spaces and short corridor components in between, providing more space for agents to avoid collisions.
EECBS. Figure 4 shows the archive of maps for EECBS. Compared to the archive of CBS, it is harder to generate challenging maps for EECBS. Map (b) and Map (c) are hard maps with typical patterns for EECBS. Map (b) contains a large chunk of empty space, but with many long corridors and one-tile entries. Map (c) contains many long corridors and one-tile entries. EECBS runs out of time on both maps. To further validate the results, we run EECBS with 200 different MAPF instances on these two maps. We get a success rate of only 11% with an average CPU runtime of 16 seconds, verifing that long corridors with one-tile entries are challenging map patterns for EECBS to solve. On the other hand, we observe that EECBS performs well in maps with more empty spaces between each long obstacle component and maps with short obstacle components, where obstacle components refer to clusters of two or more obstacles. Map (a) and Map (d) are two maps with typical patterns.
Map (a) contains many long obstacle components with one-tile entries; however, the map has two or more columns of empty space between each long obstacle component. Map (d) contains many short obstacle components with many entries in between, providing more space for agents to avoid collision. Other maps where EECBS performs well are of similar patterns or have a large chunk of empty spaces. With more entry space, it is easier for EECBS to resolve collisions. We show more maps with similar patterns in Appendix B to support our findings.
@@
PBS. Figure 5 shows the archive of maps for PBS. We empirically discover that PBS returns no solutions in many maps. Therefore, as a side discovery, we show the maps in which PBS returns no solution in at least one out of five MAPF instances during the evaluation in the archive as purple cells.
Map (c) and Map (d) are hard maps with typical patterns for PBS. Map (c) contains many long corridors with one-tile entries in between, and Map (d) contains many one-entry spaces. PBS always reports no solution on Map (c) and runs out of time on Map (d). Upon running 200 MAPF instances on Map (c) and (d), PBS could hardly solve one instance in Map (c) and always reports no solution. On Map (d), PBS only achieves a success rate of 17.5% and an average runtime of 19.1 seconds for successful instances, which validates our observation.
On the other hand, Map (a) and (b) show two typical patterns where PBS performs well. Map (a) contains long corridors but with more entry spaces in between. Map (b) contains many short obstacle components, providing more space for agents to avoid collisions. To validate our observation with long corridors, we run PBS in 200 MAPF instances on Map (a). The results indicate that PBS achieves an 87% success rate, with an average runtime of 3.4 seconds for successful cases, which indicates that PBS can effectively solve maps featuring long corridors with increased entry spaces in between.
@@
PIBT. Figure 6 shows the archive of maps of PIBT. The success rate of PIBT decreases with an increased number of obstacles and KL divergence. Similar to EECBS and CBS, PIBT performs poorly in maps with long corridors, such as Map (c). Running PIBT 200 in MAPF instances on Map (c) resulted in an average success rate of 12%, validating its poor performance. However, there are cases with large number of obstacles and low KL divergence in that PIBT has a high success rate. These maps share a common pattern: large chunks of empty spaces with few corridors. Map (a) is one such example. Running PIBT in 200 instances on Map (a) results in an average success rate of 96%. By inspecting the patterns in maps where PIBT has around a 50% success rate, we observe that the most common patterns are a combination of long corridors and one-entry spaces. Map (d) is an example. Map (b), on the other hand, has many one-entry spaces but fewer obstacles, achieving a success rate of 100%. Upon running PIBT in map (b) with 200 MAPF instances, PIBT achieves an average success rate of 95%. This result indicates that PIBT performs well when encountering one-entry spaces with fewer obstacles. Overall, PIBT can efficiently solve maps that include long corridors and one-entry spaces if given sufficient space.
LTF. Figure 7 shows the archive of maps for LTF. Similar to PIBT, LTF performs worse with more obstacles and higher KL divergence, as shown in Map (a). Similar to challenging maps for CBS, EECBS, and PIBT, Map (a) contains long corridors, which increase the chance of congestion at the entries. We use the manually designed maps with long corridors, same as the maps tested with PBS, to further evaluate the performance of LTF. The result shows that with a makespan of 512, the success rate for LTF is below 10%. With fewer obstacles, LTF performs poorly on maps with many one-entry spaces, such as Map (b) and Map (d). These two maps contain different numbers of obstacles, but they have similar patterns of one-entry spaces. Upon running LTF in 200 instances on each map, we observe an average success rate of 35%, revealing poor performance on maps with this pattern. On the other hand, LTF performs better with fewer obstacles. These maps possess large chunks of empty spaces, exemplified in Map (c).
@@ 5.1 Experiment Setup
We run EECBS with $w=1.5$. For EECBS and PBS, we use $T=20$ seconds. For PIBT and LTF, we use $M=512$, which is used to train the LTF policy (Skrynnik et al. 2024). We use the same $O_{lb}$, $O_{ub}$, $N_{e}$, $N_{eval}$, $N_{a}$ and map sizes as the one-algorithm experiments. The implementation and compute resources are specified in Section 4.1.
@@ 5.2 Results
We show the archives and representative benchmark maps in Figures 8 and 9. To better visualize the comparisons of the algorithms, we plot the differences, instead of the absolute differences, between the CPU runtime or success rate of the algorithms in each map. In Figure 8, the yellow cells represent $t_{\text{EECBS}}(\boldsymbol{x})-t_{\text{PBS}}(\boldsymbol{x})\geq 5$ seconds, and the dark blue cells represent $t_{\text{PBS}}(\boldsymbol{x})-t_{\text{EECBS}}(\boldsymbol{x})\geq 5$ seconds. In Figure 9, the yellow cells represent $SR_{\text{PIBT}}(\boldsymbol{x})-SR_{\text{LTF}}(\boldsymbol{x})\geq 30\%$, and the blue cells represent $SR_{\text{LTF}}(\boldsymbol{x})-SR_{\text{PIBT}}(\boldsymbol{x})\geq 30\%$.
@@
EECBS vs. PBS. Figure 8 shows the archive comparing EECBS and PBS. In most cases, EECBS and PBS have similar performances. However, we can identify their advantageous ranges clearly at the upper right corner and bottom left corner. Maps with more than 512 obstacles are usually too hard for both algorithms. Therefore, it is hard to generate maps that have a large difference in CPU runtime. With fewer than 400 obstacles, EECBS generally performs better than PBS in two types of maps, exemplified in Map (a) and (d). Map (a) is similar to Map (a) of Figure 4, with long but wide corridors and one-entry spaces between adjacent corridors. Map (d) shows a relatively random pattern with many one-entry spaces. To validate the comparison, we run 200 MAPF instances in Map (a) and (d). In Map (a), EECBS has a higher success rate (95.5%) than PBS (69.5%), and a lower average runtime (4.3 seconds) than PBS (12.5 seconds). Similarly, in Map (d), EECBS has a higher success rate (100%) than PBS (77.5%), and a lower average runtime (1.13 seconds) than PBS (7.4 seconds).
Map (b) and (c) are examples where PBS outperforms EECBS. Notably, while Map (c) contains narrow long corridors, both ends of the corridors are open, potentially making PBS more easily resolve conflicts. Upon validating our results on 200 MAPF instances, in Map (b), PBS has a higher success rate (96.5%) than EECBS (56.5%), and a lower average runtime (2.0 seconds) than EECBS (9.8 seconds). In Map (c), PBS has a higher success rate (43%) than EECBS (14.5%), and a lower average runtime (13.7) than EECBS (18.3 seconds).
@@
Our results serve as a more comprehensive comparison between EECBS and PBS. Notably, only one work (Chan et al. 2023) has systematically compared EECBS and PBS. They conclude that EECBS outperforms PBS in five out of six maps selected from the MAPF benchmark (Stern et al. 2019). In contrast, our results indicate that EECBS and PBS exhibit similar behavior on most maps, with each having distinct advantages on different map patterns.
@@
PIBT vs. LTF. Figure 9 shows the archive comparing PIBT and LTF. In most cases with less than 50% of obstacles, PIBT and LTF have a similar success rate. In maps with large chunks of empty space such as Map (d), both PIBT and LTF perform well. With the increase of obstacles over 50%, in general, PIBT has a 5% to 10% higher success rate on average than LTF. In most cases, these maps contain many one-entry spaces, such as Map (a). Upon running 200 instances on Map (a) with PIBT and LTF, PIBT outperforms LTF by an average success rate of 13%. The result aligns with our observations in Section 4. PIBT has better performance in maps with many one-entry spaces compared to LTF.
On the other hand, the archive also shows many cases where LTF has a higher success rate than PIBT. These cases are dispersed throughout the archive, with notable concentrations in the upper right and bottom left. Map (b) and (c) in Figure 9 are the representative maps. Upon running 200 instances on both maps, LTF has an average of 20% higher success rate on these maps compared to PIBT. By examining the patterns in these maps, we find that a slight increase in the number of empty spaces between corridors can create a significant gap in success rates between LTF and PIBT.
@@ 6 Discussion and Conclusion
When and How to Use the Proposed Framework. While developing novel MAPF algorithms, researchers can use our one-algorithm pipeline with their desired objective to find maps their algorithm has difficulty solving. By generating such maps, they can have a better understanding of the weaknesses of their algorithms. In addition, researchers can leverage our two-algorithm pipeline to compare their proposed algorithm with an existing algorithm. They can observe which kinds of maps their algorithm performs better or worse than the existing algorithm. We argue that this is a more systematic way of comparing two algorithms than comparing them on a pre-defined set of fixed benchmark maps because researchers might intentionally or unintentionally cherry-pick maps that favor their algorithm or overfit the design of their algorithms to the selected set of maps.
Nevertheless, the one-algorithm and two-algorithm pipelines are not the only realizations of our framework. Researchers can tailor the framework for alternative purposes by designing the objectives and diversity measures. For example, researchers can minimize, instead of maximize, the CPU runtime of CBS, EECBS, and PBS with a larger number of agents to generate a diverse collection of maps that are easy, instead of hard, for these algorithms. While we present some easy maps for each selected algorithm in Section 4, they are by no means an exhaustive set. If easy maps are of interest, researchers should use alternative objectives, such as minimizing the CPU runtime or maximizing the success rate. If researchers are interested in both hard and easy maps, using the CPU runtime or regularized success rate as diversity measures is a better choice.
@@
Computational Cost. Our proposed framework relies on QD algorithms, which require a massive number of evaluations in the generated benchmark maps to compute the objective and diversity measure values. Specifically, with our one-algorithm experiment setup of $N_{eval}=10,000$ evaluations, each with $N_{e}=5$ runs in different MAPF instances, it takes 6 to 7 hours to generate benchmark maps for PIBT and LTF, 10 hours for PBS, and more than 12 hours for CBS and EECBS. The runtime is measured in machine (1) specified in Section 4.1. Future works can focus on reducing the computational cost of QD algorithms to make the pipeline more efficient. For example, prior works (Zhang et al. 2022; Bhatt et al. 2022, 2023) have explored training a data-driven surrogate model to assist the QD search. Meanwhile, we will release our most representative generated benchmark maps online. If the researchers’ new algorithm is built on one of the algorithms we have tested, they can directly evaluate new algorithms on our benchmark maps.
Acknowledgments
The research at Carnegie Mellon University was supported by the National Science Foundation (NSF) under grant number #2328671 and a gift from Amazon. The work used Bridge-2 at Pittsburgh Supercomputing Center (PSC) through allocation CIS220115 from the Advanced Cyberinfrastructure Coordination Ecosystem: Services & Support (ACCESS) program, which is supported by NSF under grant numbers #2138259, #2138286, #2138307, #2137603, and #2138296. The views and conclusions contained in this document are those of the authors and should not be interpreted as representing the official policies, either expressed or implied, of the sponsoring organizations, agencies, or the U.S. government. We also thank Daniel Harabor, Peter J. Stuckey, Nathan Sturtevant, as well as reviewers at AAAI 2025 for their valuable feedback on this paper.
@@ Appendix A Choosing Diversity Measures
In this section, we justify our choice of diversity measures by setting up the archive with every two pairs of competitive diversity measures and exploring their relationship and behavior. We show the following experiments: (1) we compare the number of obstacles and standard deviation of Betweenness Centrality (std of BC)  (Ewing et al. 2022), the number of obstacles and $\lambda_{2}$ (Ren et al. 2024), and show that the number of obstacles could quantify the general hardness of the generated maps better than std of BC and $\lambda_{2}$. (2) We compare the KL divergence of the tile pattern distribution and map entropy, the KL divergence of the tile pattern distribution and the KL divergence of the WL graph feature, and show that the KL divergence of the tile pattern distribution can generate more complex patterns than map entropy and have less computational cost than the KL divergence of WL graph feature.
Figure 10: Archive of PBS with the number of obstacles and std of BC as diversity measures.

### v4

@@ 4 Benchmarking One Algorithm
For one-algorithm experiments, we intend to generate a diverse set of difficult benchmark maps targeted for CBS, EECBS, PBS, PIBT, and LTF.
@@
Figure 2: Examples of one-entry space and one-tile entry.
@@ 4.1 Experiment Setup
Hyperparameters. Defining $w$ as the suboptimality bound away from the optimal solution, we use $w=1.5$ for EECBS. For CBS, we use the implementation of EECBS with $w=1$. We set $T=20$ seconds for CBS, EECBS, and PBS, $M=1000$ for PIBT, and $M=512$ for LTF. For all algorithms, we fix the map size as 32 $\times$ 32 and the number of agents at 150. The pre-defined range for the number of obstacles is $[O_{lb}=307,O_{ub}=717]$, which corresponds to 30% to 70% of the map size. For each evaluation, we run the given MAPF algorithm in $N_{e}=5$ different instances. For all algorithms, we evaluated $N_{eval}=10,000$ maps. For CBS, EECBS, and PBS, we run evaluations with $N_{a}=50$ agents, and for PIBT and LTF, we use $N_{a}=150$. We use more agents for PIBT and LTF because they empirically have higher success rates in congested MAPF instances. We include implementation details and compute resources in Appendix B.
@@ 4.2 Results
Generated Maps. We first define one-entry space and one-tile entry (shown in Figure 2) as features of the benchmark maps. A one-entry space represents the empty space surrounded by three obstacles. A one-tile entry represents a single empty space separating two corridors. In Figures 3, 4 and 5, the yellow and dark blue cells indicate maps with high and low CPU runtime, respectively. In Figures 6 and 7, the yellow and dark blue cells indicate maps with low and high success rates, respectively. The colored tiles in the maps display 10 out of 150 pairs of start and goal locations. While we are primarily interested in maps that are challenging for the algorithms, we also present easy maps as the by-products of the experiments. We summarize the typical patterns of the maps and show more maps with similar patterns in Appendix C.
@@
CBS. Figure 3 shows the archive of maps for CBS. It is easy to generate challenging maps for CBS. Map (a), Map (c), and Map (d) show the most typical hard maps for CBS. Map (a) and Map (c) contain many long corridors, and Map (d) contains many one-entry spaces with long corridors. Meanwhile, we also find that CBS can easily solve maps with fewer obstacles and a large chunk of empty spaces, such as Map (b). It provides more space for agents to avoid collisions.
EECBS. Figure 4 shows the generated maps for EECBS. Compared to CBS, it is harder to generate challenging maps for EECBS. Map (b) and Map (c) are typical hard maps for EECBS. Map (b) contains a large chunk of empty space, but with many long corridors and one-tile entries. Map (c) contains many long corridors and one-tile entries. EECBS runs out of time on both maps. On the other hand, EECBS performs well in maps with more empty spaces between each long obstacle component (Map (a)) and maps with short obstacle components (Map(d)), where obstacle components refer to connected components of two or more obstacles.
@@
PBS. Figure 5 shows the archive of maps for PBS. We discover that PBS returns no solutions in many maps. Therefore, as a side discovery, we show the maps in which PBS returns no solution in at least one out of five MAPF instances during the evaluation in the archive as orange cells. Map (c) and Map (d) are typical hard maps for PBS. Map (c) contains many long corridors with one-tile entries in between, and Map (d) contains many one-entry spaces. PBS always reports no solution on Map (c) and runs out of time on Map (d). On the other hand, Map (a) and (b) show two typical easy maps for PBS. Map (a) contains long corridors but with more entry spaces in between. Map (b) contains many short obstacle components, providing more space for agents to avoid collisions.
@@
PIBT. Figure 6 shows the archive of maps of PIBT. The success rate of PIBT decreases with an increased number of obstacles and KL divergence. Similar to EECBS and CBS, PIBT performs poorly in maps with long corridors, such as Map (c). With a similar number of obstacles and smaller KL divergence, we can find easy maps for PIBT. They have large chunks of empty spaces with few corridors. Map (a) is one such example. In the middle of the archive, we observe maps with success rates of around 50%. They usually have long corridors and one-entry spaces. Map (d) is an example. In the bottom left corner, we find Map (b), which has many one-entry spaces but fewer obstacles, achieving a success rate of 100%. Overall, PIBT can efficiently solve maps that include long corridors and one-entry spaces if given sufficient space.
@@
LTF. Figure 7 shows the archive of maps for LTF. Similar to PIBT, LTF performs worse with more obstacles and higher KL divergence, as shown in Map (a). Similar to challenging maps for CBS, EECBS, and PIBT, Map (a) contains long corridors. With fewer obstacles, LTF performs poorly on maps with many one-entry spaces, such as Map (b) and Map (d). These two maps contain different numbers of obstacles, but they have similar patterns of one-entry spaces. On the other hand, LTF performs better with fewer obstacles. These maps possess large chunks of empty spaces, such as Map (c).
Validation Results. We only evaluate each map by running $N_{e}=5$ MAPF instances during the QD search. To further validate our observations about the hardness of the maps, we additionally present validation results in Table 1 of Appendix D by running 200 MAPF instances in the selected maps. The results validate the hardness of the maps.
@@ 5.1 Experiment Setup
We run EECBS with $w=1.5$. For EECBS and PBS, we use $T=20$ seconds. For PIBT and LTF, we use $M=512$, which is used to train the LTF policy Skrynnik et al. (2024). We use the same $O_{lb}$, $O_{ub}$, $N_{e}$, $N_{eval}$, $N_{a}$, and map sizes as the one-algorithm experiments.
@@ 5.2 Results
Generated Maps. We show the archives and representative benchmark maps in Figures 8 and 9. To better visualize the comparisons of the algorithms, we plot the differences, instead of the absolute differences, between the CPU runtime or success rate of the algorithms. In Figure 8, the yellow and dark blue cells represent EECBS having lower and higher CPU runtime than PBS, respectively. In Figure 9, the yellow and blue cells represent PIBT having higher and lower success rates than LTF, respectively.
@@
EECBS vs. PBS. Figure 8 shows the archive comparing EECBS and PBS. In most cases, EECBS and PBS have similar performances. Maps with more than 512 obstacles are usually too hard for both algorithms. With fewer than 400 obstacles, however, EECBS generally performs better than PBS in two types of maps, exemplified in Map (a) and (d). Map (a) is similar to Map (a) of Figure 4, with long but wide corridors and one-entry spaces between adjacent corridors. Map (d) shows a random pattern with many one-entry spaces. On the other hand, Map (b) and (c) are examples where PBS outperforms EECBS. Notably, while Map (c) contains narrow long corridors, both ends of the corridors are open, potentially making PBS resolve conflicts easier.
@@
Our results serve as a more comprehensive comparison between EECBS and PBS. Notably, only one work Chan et al. (2023) has systematically compared EECBS and PBS. They conclude that EECBS outperforms PBS in five out of six maps selected from the MAPF benchmark Stern et al. (2019). In contrast, our results indicate that EECBS and PBS exhibit similar behavior on most maps, with each having distinct advantages on different map patterns.
@@
PIBT vs. LTF. Figure 9 shows the archive comparing PIBT and LTF. In most cases with less than 50% of obstacles, PIBT and LTF have a similar success rate. In maps with large chunks of empty space such as Map (d), both PIBT and LTF perform well. With more than 50% of obstacles, PIBT generally has a 5% to 10% higher success rate on average than LTF. In most cases, these maps contain many one-entry spaces, such as Map (a). On the other hand, the maps in which LTF outperforms PIBT are concentrated in the upper right and bottom left, such as Map (b) and (c).
@@
Validation Results. Similar to Section 4, we run 200 MAPF instances in selected maps and show the results in Table 2 of Appendix D to validate the hardness of the generated maps.
@@ 6 Discussion and Conclusion
When and How to Use the Proposed Framework. While developing novel MAPF algorithms, researchers can use our one-algorithm pipeline to find maps their algorithm has difficulty solving. By generating such maps, they can have a better understanding of the weaknesses of their algorithms. Researchers can also leverage our two-algorithm pipeline to compare their proposed algorithm with an existing algorithm. They can observe which kinds of maps their algorithm performs better or worse than the existing algorithm. We argue that this is a more systematic way of comparing two algorithms than comparing on a fixed set of benchmark maps because researchers might intentionally or unintentionally cherry-pick maps that favor their algorithm or overfit the design of their algorithms to the selected set of maps.
Nevertheless, the pipelines shown in the paper are not the only realizations of our framework. Researchers can tailor the framework for alternative purposes by designing the objectives and diversity measures. For example, researchers can minimize, instead of maximize, the CPU runtime of CBS, EECBS, and PBS with a larger number of agents to generate maps that are easy, instead of hard, for these algorithms. While we present some easy maps for each selected algorithm in Section 4, they are by no means an exhaustive set. If easy maps are of interest, researchers should use alternative objectives, such as minimizing the CPU runtime or maximizing the success rate. If researchers are interested in both hard and easy maps, using the CPU runtime or regularized success rate as diversity measures is a better choice.
@@
Computational Cost. Our proposed framework relies on QD algorithms, which require a massive number of evaluations in the generated benchmark maps. Specifically, with our one-algorithm experiment setup, it takes 6 to 7 hours to generate benchmark maps for PIBT and LTF, 10 hours for PBS, and more than 12 hours for CBS and EECBS. Future works can focus on reducing the computational cost of QD algorithms by using surrogate models Zhang et al. (2022). Meanwhile, we will release our most representative generated benchmark maps online. If the researchers’ new algorithm is built on one of the algorithms we have tested, they can directly evaluate new algorithms on our benchmark maps.
@@ Appendix A Choosing Diversity Measures
In this section, we justify our choice of diversity measures by setting up the archive with every two pairs of competitive diversity measures and exploring their relationship and behavior. We show the following experiments: (1) we compare the number of obstacles and standard deviation of Betweenness Centrality (std of BC)  Ewing et al. [2022], the number of obstacles and $\lambda_{2}$ Ren et al. [2024], and show that the number of obstacles could quantify the general hardness of the generated maps better than std of BC and $\lambda_{2}$. (2) We compare the KL divergence of the tile pattern distribution and map entropy, the KL divergence of the tile pattern distribution and the KL divergence of the WL graph feature, and show that the KL divergence of the tile pattern distribution can generate more complex patterns than map entropy and has less computational cost than the KL divergence of WL graph feature.
Figure 10: Scatter plot of about 4370 generated benchmark maps.